# Autoformalization in the Era of Large Language Models: A Survey

### v1

@@ 3.3 Therom Proof
While autoformalization emphasizes the transformation of informal statements into formal representations, its practical utility is ultimately grounded in the ability to construct complete and verifiable proofs. As such, recent advances in large language model–driven theorem proving not only complement autoformalization but also expand its scope—by generating formal derivations that validate the translated statements. These methods can be broadly categorized by their granularity: some focus on the step-by-step generation of tactics, enabling precise control and local reasoning, while others aim for holistic, end-to-end proof generation. We begin by reviewing the former, which centers around tactic-level reasoning and search, and plays a foundational role in training models to navigate the formal proof space effectively.
@@ 3.3.1 Tactic Step Proving Methods
Recently, many works have been devoted to enhancing the local reasoning ability of automated theorem proving using tactic steps. For example, Xin et al. (2025) proposes a scalable expert iteration framework based on breadth-first tree search. In each iteration, the framework filters the question set, and combines Direct Preference Optimization and length normalization strategies to optimize the language model. It makes it preferentially explore promising proof paths. Bfs-prover achieves a 72.95% pass rate on the Lean4 MiniF2F benchmark, showing that simple BFS search can be competitive under appropriately scaled conditions.
@@ 4.2.1 High School Level
High school math problems mainly focus on high school math competition problems, such as IMO, AIMC and other math competitions. Problems usually have a clear idea, a short proof process, and a standardized logical structure, and generally involve basic algebraic, geometric, and preliminary number theory propositions. Zheng et al. (2022) proposes a representative data set miniF2F problems are mostly taken from high school Olympiad math problems, with clear structure and limited steps.
@@ 4.2.2 Undergraduate Level
Undergraduate-level math problems are mostly from classical math textbooks, covering the fields of real analysis, linear algebra, abstract algebra, topology, etc. The representative data set is ProofNet presented in Azerbayev et al. (2023), which contains 371 examples, each of which is composed of formal theorems in Lean 3, natural language theorem statements and proofs. It covers real analysis, complex analysis, linear algebra, abstract algebra and topology. The FormL4 dataset presented in Lu et al. (2024b) extracts theorems from Mathlib4 and generates corresponding informal descriptions through a large language model, reflecting the step-by-step progression of mathematical theorems from foundation to complex proofs in university. The problem involves a higher level of abstraction, the proof process is long and rigorous, and the model needs to understand high-level mathematical concepts.
@@ 4.2.3 Graduate Level
Problems at the graduate level and beyond are usually derived from the latest academic papers or advanced mathematical monographs, and involve a higher degree of theorems background, proof strategies, and conceptual abstraction than at the undergraduate level. Patel et al. (2024) proposes a representative dataset named arXiv2Formal, sampled from mathematical papers on arXiv, and contains 50 research-level theorems, focusing on ”unlinked formalization”, i.e. when the necessary background definition is missing in the formalization process, and the model needs to be ”unlinked formalization”, i.e. in the absence of a full context, the model needs to be ”unlinked formalization”. Complete the connection between definitions, lemmas and theorems. Models need to deal with highly abstract mathematical concepts, understand complex proof strategies, and reason and link in the absence of full context; The Herald dataset proposed by Gao et al. (2025) provides natural language notes and formal proof pairs generated by Mathlib4 in Lean 4, aiming to provide fine-grained corpus of advanced mathematical proofs for automatic formal systems. It is mainly for graduate students and above in mathematical research, and contains high-level mathematical languages and proof strategies. It is suitable for training models to deal with complex mathematical proof tasks.
@@ 4.3.1 Procedural Abstraction in High School Mathematics
According to Advanced Mathematical Thinking, the understanding of functions at the high school level typically remains at the stage of procedural abstraction—characterized by reliance on techniques rather than formal axiomatic reasoning. High school mathematics is relatively low in abstraction, with content tightly organized around concrete examples and intuitive interpretations. For example, finding the zeros of a quadratic function usually involves applying procedural techniques such as computing the discriminant or analyzing the graph of the function. Functional properties like monotonicity or evenness are verified through direct algebraic manipulation, with strong dependence on numerical values rather than logical formalism. For instance, to demonstrate the monotonicity of $f(x)=e^{x}$, one computes its derivative—a method grounded in concrete computation rather than abstract reasoning.
@@ 4.3.3 Structural Abstraction in Graduate Mathematics
5 Connecting LLMs and Formal Verification
Autoformalization plays a crucial role in the formal verification of large model outputs. Although LLMs have powerful performance, they also suffer from serious ”hallucination”, which makes the output of LLMs cannot be completely trusted. Therefore, the formal verification of large model output has become a hot research direction. However, in order to get fully trusted output, the LLMs need to make completely correct reasoning, which is essentially a mathematical problem, and can use various formal verification methods, such as first-order logic, higher-order logic, formal languages, etc. No matter what kind of formal verification method is used, the common problem is how to autoformalize the output into the corresponding ”rules” formulated by the method. In the direction of promoting the further development of LLMs, autoformalization is a bridge between LLMs and formal verification technology. Making the two advantages complement each other and making the LLMs generate fully credible output becomes the goal of the next stage of research.
@@ 5.3 LLM output verification
Song et al. (2025) explores a framework in which a model validates its own output. Based on this verification, filter or reweight the data and extract the filtered data. And a comprehensive, modular and controlled LLM self-improvement study was initiated. This article provides a mathematical formula for self-improvement, mainly controlled by a formal generative validation gap (GV-Gap). The SymGen method proposed in Hennigen et al. (2023) uses ”symbolic references” to establish associations between natural language texts and structured data. Symbolic references provide clear data source references for autoformalization by embedding reference identifiers for specific data fields (such as filename) in the text. This reference mechanism is similar to model annotation or marking in formal methods, which can help formal tools quickly locate the data elements on which the text content depends, thereby achieving more efficient data-driven formal transformation. It not only helps ensure the accuracy and traceability of the content in the text generation stage, but also plays a key role in the subsequent formal verification and maintenance process.
6 Foundations of Progress: Techniques and Corpora
  • Autoformalized proof: Using the reasoning ability of LLMs, the given natural language mathematical problem statement is automatically generated into verifiable formal proof.
  • Data construction: Through large-scale synthesis and generation, enhance or filter data and other methods to break through the bottleneck of data scarcity and improve the performance of the model.
  • Evalute Autoformalization: Compilation of formal proof assistans does not mean that autoformalization is completely correct. In order to better verify the correctness of autoformalization, new methods or performance metrics are proposed.
  • Retrieval Augmented: Given the current state of the proof, enhance the ability to retrieve potential premises and establish reflection between them. When proving a theorem, the model generates multiple candidate strategies at each step, and the proof can be found by the most priority search algorithm, thus establishing the correct proof.
  • DSP: Through a LLM, the informal proof is mapped to the formal proof sketch, and the mapped sketch is used to help the automatic prover search for subproblems that are simpler than the complete problem to complete the proof.
Table 2 presents a comprehensive overview of datasets for theorem proving and autoformalization. The datasets are categorized by mathematical difficulty level (e.g. Olympiad-level, undergraduate mathematics), size (number of problems/theorems), primary task, and brief descriptions. Understanding the mathematical difficulty levels of these datasets is critical, as it directly impacts model evaluation: lower-level problems (e.g. grade school math) test basic reasoning, while advanced datasets (e.g. research-level formalizations) assess a model’s capacity to handle abstraction and complex dependencies. The main task is to generalize the category of tasks supported by the dataset, including theorem proving, autoformalization, etc. The Task column of the table is explained as follows:
  • Theorem Proving: Generating or verifying machine-checkable formal proofs from given statements, often using predefined tactics or proof steps .
  • Autoformalization: Translating informal mathematical statements into formal code while preserving semantic equivalence.
  • Fine-tuning: Adapting pre-trained models to specialized domains (e.g., competition math) using smaller, task-specific datasets like Math (12.5k problems).
  • Pre-training: Training foundational models on large-scale corpora to learn general mathematical reasoning patterns.
  • Question Answering: Solving math problems and generating step-by-step natural language or formalized answers.

### v2

@@ 4.3 Therom Proof
While autoformalization emphasizes the transformation of informal statements into formal representations, its practical utility is ultimately grounded in the ability to construct complete and verifiable proofs. As such, recent advances in large language model–driven theorem proving not only complement autoformalization but also expand its scope—by generating formal derivations that validate the translated statements. Azerbayev et al. (2024) suggests LLEMMA, 7B and 34B parametric language models specifically designed for mathematics, which leverage the few-shot learning capabilities of LLM to achieve good performance on several benchmarks. Xin et al. (2024a) released the base model DeepSeekMath with 7B parameter sizes achieving comparable performance to Minerva540B, which also shows that the number of parameters is not the only key factor in mathematical reasoning ability, and strong performance can also be obtained in smaller models. These methods can be broadly categorized by their granularity: some focus on the step-by-step generation of tactics, enabling precise control and local reasoning, while others aim for holistic, end-to-end proof generation. We begin by reviewing the former, which centers around tactic-level reasoning and search, and plays a foundational role in training models to navigate the formal proof space effectively.
@@ 4.3.1 Tactic Step Proving Methods
Recently, many works have been devoted to enhancing the local reasoning ability of automated theorem proving using tactic steps. For example, Xin et al. (2025b) proposes a scalable expert iteration framework based on breadth-first tree search. In each iteration, the framework filters the question set, and combines Direct Preference Optimization and length normalization strategies to optimize the language model. It makes it preferentially explore promising proof paths. Bfs-prover achieves a 72.95% pass rate on the Lean4 MiniF2F benchmark, showing that simple BFS search can be competitive under appropriately scaled conditions.
@@ 5.2.1 High School Level
High school math problems mainly focus on high school math competition problems, such as IMO, AIMC and other math competitions. Problems usually have a clear idea, a short proof process, and a standardized logical structure, and generally involve basic algebraic, geometric, and preliminary number theory propositions. Zheng et al. proposes a representative data set miniF2F problems are mostly taken from high school Olympiad math problems, with clear structure and limited steps.
@@ 5.2.2 Undergraduate Level
Undergraduate-level math problems are mostly from classical math textbooks, covering the fields of real analysis, linear algebra, abstract algebra, topology, etc. The representative data set is ProofNet presented in Azerbayev et al. (2023), which contains 371 examples, each of which is composed of formal theorems in Lean 3, natural language theorem statements and proofs. It covers real analysis, complex analysis, linear algebra, abstract algebra and topology. Yu et al. (2025) and Liu et al. (2025a) include theorems from both high school Olympic competitions and undergraduate levels. Yu et al. (2025) covers a variety of fields, while Liu et al. (2025a) and Xiong et al. (2025) both focus on building benchmarks for combinatorial math problems. The FormL4 dataset presented in Lu et al. (2024b) extracts theorems from Mathlib4 and generates corresponding informal descriptions through a large language model, reflecting the step-by-step progression of mathematical theorems from foundation to complex proofs in university. The problem involves a higher level of abstraction, the proof process is long and rigorous, and the model needs to understand high-level mathematical concepts.
@@ 5.2.3 Graduate Level
Problems at the graduate level and beyond are usually derived from the latest academic papers or advanced mathematical monographs, and involve a higher degree of theorems background, proof strategies, and conceptual abstraction than at the undergraduate level. Patel et al. proposes a representative dataset named arXiv2Formal, sampled from mathematical papers on arXiv, and contains 50 research-level theorems, focusing on “unlinked formalization”, i.e. when the necessary background definition is missing in the formalization process, and the model needs to be “unlinked formalization”, i.e. in the absence of a full context, the model needs to be “unlinked formalization”. Complete the connection between definitions, lemmas and theorems. Models need to deal with highly abstract mathematical concepts, understand complex proof strategies, and reason and link in the absence of full context; The Herald dataset Gao et al. (2025) provides natural language notes and formal proof pairs generated by Mathlib4 in Lean 4, aiming to provide fine-grained corpus of advanced mathematical proofs for automatic formal systems. It is mainly for graduate students and above in mathematical research, and contains high-level mathematical languages and proof strategies. It is suitable for training models to deal with complex mathematical proof tasks.
@@ 5.3.1 Procedural Abstraction in High School Mathematics
According to Tall (1991), the understanding of functions at the high school level typically remains at the stage of procedural abstraction—characterized by reliance on techniques rather than formal axiomatic reasoning. High school mathematics is relatively low in abstraction, with content tightly organized around concrete examples and intuitive interpretations. For example, finding the zeros of a quadratic function usually involves applying procedural techniques such as computing the discriminant or analyzing the graph of the function. Functional properties like monotonicity or evenness are verified through direct algebraic manipulation, with strong dependence on numerical values rather than logical formalism. For instance, to demonstrate the monotonicity of $f(x)=e^{x}$, one computes its derivative—a method grounded in concrete computation rather than abstract reasoning.
@@ 5.3.3 Structural Abstraction in Graduate Mathematics
6 Connecting LLM Outputs and Formal Verification via Autoformalization
Autoformalization plays a crucial role in applying formal verification methods to large language model (LLM) outputs. Despite their powerful capabilities, LLMs often suffer from significant hallucinations, undermining the reliability and trustworthiness of their generated results. Consequently, verifying the correctness and trustworthiness of LLM-generated outputs has become a pressing research topic. Achieving fully trustworthy outputs requires rigorous verification of the underlying reasoning processes, which essentially translates into mathematical and logical verification tasks. Such tasks can leverage various formal verification frameworks, including first-order logic, higher-order logic, and formal specification languages. Regardless of the chosen formal verification approach, a common challenge is accurately translating—or autoformalizing—the informal outputs produced by LLMs into formal representations suitable for verification. Autoformalization thus serves as a critical bridge connecting LLM reasoning with formal verification techniques.
@@ 6.3 LLM output verification
Song et al. explores a framework in which a model validates its own output. Based on this verification, filter or reweight the data and extract the filtered data. And a comprehensive, modular and controlled LLM self-improvement study was initiated. This article provides a mathematical formula for self-improvement, mainly controlled by a formal generative validation gap (GV-Gap). The SymGen method Hennigen et al. (2023) uses “symbolic references” to establish associations between natural language texts and structured data. Symbolic references provide clear data source references for autoformalization by embedding reference identifiers for specific data fields (such as {filename}) in the text. This reference mechanism is similar to model annotation or marking in formal methods, which can help formal tools quickly locate the data elements on which the text content depends, thereby achieving more efficient data-driven formal transformation. It not only helps ensure the accuracy and traceability of the content in the text generation stage, but also plays a key role in the subsequent formal verification and maintenance process.